OpenStudy (anonymous):

Solve the systems of equations x + 3y = 32 -3x + 2y = 3 and x + y = 5 x - y = -9

OpenStudy (anonymous):

x=32-3y pluggin this in the 2nd eq. -3(32-3y)+2y=3 -96+9y+2y=3 11y=3+96 y=99/11=9 pluggin y=9 in any of the two eq we get x=5

OpenStudy (anonymous):

y=9 x=5

OpenStudy (anonymous):

x = -2 y=7

OpenStudy (anonymous):

proceed the same way for the second Q... try it..its vry easy

OpenStudy (anonymous):

Thanks so much for the help!

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

Ok i did the second problem and got y = -5 here is my work can you tell me where i went wrong? x + y = 5 x = 5- y x - y = -9 (5-y) –y = -9 -5y-y = -9 -4y = -9 y = -5

OpenStudy (anonymous):

(5-y) –y = -9 check it from here (5-y) –y = -9 5-y-y=-9 the two -y add to give -2y 5-2y=9 2y=5-9 y=-4/2=-2 did you understand?

OpenStudy (anonymous):

Ah I see now... i'm really terrible at math, don't know why it's so hard for me to grasp. Thanks again!

OpenStudy (anonymous):

keep practicing.. thats the only way to success=)